# Quality Tagging, and Management of Synthetic Data

### v2

@@ 1 Introduction
Figure 1: High-level SyGra architecture.
2 Related Work
Recent years have seen rapid progress in the development of synthetic data generation frameworks and instruction-tuning toolkits, with each system making distinct trade-offs across orchestration, extensibility, code abstraction, and multimodal support. Table 1 summarizes core capabilities across representative frameworks like Distilabel [^1] https://distilabel.argilla.io/latest/, SDG [^2] https://github.com/argilla-io/synthetic-data-generator, Curator [^3] https://github.com/bespokelabsai/curator, and Synthetic Data Kit [^4] https://github.com/meta-llama/synthetic-data-kit.
3 SyGra Framework
@@ 3.1 System Architecture
    Graph Construction: YAML-defined DAG of nodes (LLM calls, transformations) with conditional edges and pre/post hooks, compiled via LangGraph.
@@
    Execution Engine: Asynchronous runtime coordinating local Python steps and remote inference (HTTP, OpenAI, Mistral) across VLLM, TGI, OLLAMA and Azure backends, with built-in retries and failure tracing.
@@ 3.3 Key Features
    Multimodal Support: SyGra extends beyond text-only workflows by natively handling audio and image inputs alongside text. Through unified I/O adapters, it transparently loads local or remote media in various formats, encodes them as base64 data URLs for LLM API compatibility, and supports multiple media fields per record. This enables workflows for tasks such as speech recognition, audio classification, document analysis, and visual QA. Round-tripping ensures outputs can be saved back into HuggingFace datasets in their original formats for reproducibility and downstream use.
@@ 5 Availability
SyGra is released as an open-source Python package and framework on PyPI555pip package: https://pypi.org/project/sygra/ (pip install sygra) and GitHub [^6] https://github.com/ServiceNow/SyGra.
@@ 6 Conclusion
Current limitations include text-only outputs (with multimodal output support planned), independent node operation without cross-sample reasoning, and basic agent support. Future work will address richer agent coordination, stateful nodes, real-time graph editing, and uncertainty modeling.
@@ A.1.1 Input Sources
Sample YAML Configuration:
Figure 7: An example configuration using a HuggingFace dataset as source and applying field renaming transformation is shown below.
This configuration illustrated in Figure 7 represents:
@@ A.1.2 Transformations
Figure 8(a) shows a sample usage where the fields page, llm_extract, and type are renamed to id, text, and text_format, respectively.
(a) Example usage of RenameFieldsTransform in YAML configuration. This renames selected fields to align with graph input expectations.
(b) CombineRecords configuration: aggregates two records with shift and join logic.
(c) SkipRecords configuration: skips first and last 10 records using count-based range.
(d) Minimal YAML configuration demonstrating data-less mode with only a sink field.
Figure 8: Examples of YAML configurations used for field renaming, record combination/skipping, and minimal data sink setup.
@@
This transformation combines multiple records to form richer contextual input. It can skip from the beginning or end of the dataset, define how many records to combine, and how to shift the combination window. As shown in Figure 8(b), the configuration merges two records, joining multiple fields with newline delimiters or preserving the first record’s values.
@@
Figure 8(c) presents a simpler configuration to exclude records from the dataset, either from the start or end. This is especially useful for filtering noisy, incomplete, or structurally incompatible entries prior to processing.
@@ A.1.3 Data Less Mode
Figure 8(d) shows a minimal YAML configuration that defines only an output sink.

### v3

@@ 1 Introduction
2 Related Work
Recent years have seen rapid progress in the development of synthetic data generation frameworks and instruction-tuning toolkits, with each system making distinct trade-offs across orchestration, extensibility, code abstraction, and multimodal support. Table 1 summarizes core capabilities across representative frameworks like Distilabel[3], SDG[2], Curator[4], and Synthetic Data Kit[1].
Figure 1: High-level SyGra architecture.
3 SyGra Framework
@@ 3.1 System Architecture
    Notably, SyGra provides support for ServiceNow as a data source and sink, enabling seamless integration with ServiceNow instances [16].
@@
    Graph Construction: YAML-defined DAG of nodes (LLM calls, transformations) with conditional edges and pre/post hooks, compiled via LangGraph [10]
@@
    Execution Engine: Asynchronous runtime coordinating local Python steps and remote inference (HTTP, OpenAI, Mistral) across VLLM [9], TGI [6], OLLAMA [12], and Azure/OpenAI backends [13], with built-in retries and failure tracing.
@@ 3.3 Key Features
    Multimodal Support: SyGra extends beyond text-only workflows by natively handling audio and image inputs alongside text. Through unified I/O adapters, it transparently loads local or remote media in various formats, encodes them as base64 data URLs for LLM API compatibility, and supports multiple media fields per record. This enables workflows for tasks such as speech recognition, audio classification, document analysis, and visual QA. Round-tripping ensures outputs can be saved back into HuggingFace datasets in their original formats for reproducibility and downstream use. Additionally, SyGra supports multimodal outputs—including generated images and audio—when using GPT-based endpoints (OpenAI or Azure OpenAI), enabling end-to-end multimodal generation for OpenAI endpoints.
@@
    Metadata Tracking: SyGra includes an automatic metadata tracking system that captures comprehensive execution metrics without requiring any code changes. The system provides real-time cost tracking for multiple LLM providers (OpenAI, Azure OpenAI, Anthropic Claude on AWS Bedrock, vLLM), detailed token usage statistics, and multi-level performance monitoring at aggregate, model, and node granularities. Metrics include latency percentiles (p50, p95, p99), throughput measurements, retry and failure rates, and response code distributions. Output and metadata files share synchronized timestamps for easy correlation, and the system captures execution context including git commit information and dataset versioning for full reproducibility.
    from sygra.metadata.metadata_collector import get_metadata_collector
    collector = get_metadata_collector()
    metadata = collector.get_metadata_summary()
    # Access aggregate metrics
    stats = metadata[’aggregate_statistics’]
    print(f"Total cost: \${stats[’cost’][’total_cost_usd’]:.4f}")
    print(f"Total requests: {stats[’requests’][’total_requests’]}")
    print(f"Models used: {list(metadata[’models’].keys())}")
    print(f"Nodes executed: {list(metadata[’nodes’].keys())}")
  8. 8.
@@
4 Dual-Stage Quality Tagging
Quality control is central to synthetic data generation. SyGra implements a two-stage mechanism balancing efficiency and accuracy: fast heuristic filtering eliminates obvious low-quality samples, followed by targeted LLM-based evaluation for samples passing initial checks. This section details both stages, metadata schema, and integration with training pipelines.
4.1 Stage 1: Heuristic-Based Filtering
The first stage applies eight checks. Implementation uses thread pool with $N_{cpu}$ workers.
1. Conversation Pretokenization: Applies model-specific chat template (fetched from HuggingFace or custom) to validate format compatibility. Computes token count using tokenizer.
2. Language Detection: Uses fastText [7] for detection (99.5% accuracy on XLM-R benchmark). Concatenates all turns, detects language, computes confidence. Rejects if detected language not in target set (default: {English}) or confidence << threshold (default: 0.90). Useful for: filtering code-switched data, enforcing monolingual datasets, handling web-scraped data with mixed languages.
3. Conversation Length Check: Counts turns and validates range. Rejects if: (a) turns << min_turns (default: 2, requires at least one exchange), or (b) turns >> max_turns (default: 20, filters extremely long dialogues that are often errors or edge cases). Prevents: empty conversations, infinitely long generations (from LLM loops), single-turn data in multi-turn pipelines.
4. Metadata Tagging: Extracts structural features for downstream filtering and analysis: turn count, average turn length (chars), role distribution (% assistant vs. user), special token usage (code blocks, math symbols, citations). Stored in metadata but does not reject samples. Used later for: stratified sampling by length, balancing role distributions, identifying domain-specific data.
5. Lexical Diversity: Computes Type-Token Ratio (TTR) and Measure of Textual Lexical Diversity (MTLD) [11]:
TTR $=\frac{|\text{unique tokens}|}{|\text{total tokens}|}$ (1)
MTLD $=\frac{1}{k}\sum_{i=1}^{k}\ell_{i}\quad\text{where }\ell_{i}\text{ is length to factor $<$0.72}$ (2)
Rejects if TTR << 0.30 (highly repetitive text, e.g., "I like cats. I like dogs. I like birds."). MTLD captures diversity over longer spans, complementing TTR’s sensitivity to sample length. Filters: LLM generation loops (same phrases repeated), template-based spam, low-effort responses.
6. Perplexity Scoring: Uses domain-specific language model (default: GPT-2 small for speed) to compute perplexity:
$\text{PPL}(x)=\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\log P(x_{i}|x_{<i})\right)$ (3)
Rejects if PPL >> threshold (default: 1000, indicating low fluency). Catches: grammatically malformed text, gibberish, foreign languages not caught by detector, heavily code-mixed data. Trade-off: requires GPU for fast scoring (CPU: 50 samples/sec, GPU: 500 samples/sec). Optional—users can disable if throughput is critical.
7. Reward Modeling (Optional): Applies pre-trained reward model to predict human preference score. Rejects if score << threshold (default: 0.3). Provides semantic quality signal beyond surface features. Trade-off: expensive (requires GPU, 100-200ms per sample). Recommended for high-value datasets where cost is acceptable.
8. Data Characteristics: Computes domain-specific features used for routing and analysis: presence of code blocks, mathematical symbols, citations, tables, lists. Determines conversation category (generic, math, code, QA). Stored in metadata for Stage 2 routing. No rejection—purely informational.
4.2 Stage 2: LLM-Based Categorical Evaluation
Samples after Stage 1 undergo category-specific LLM evaluation.
Category Classification: Based on Data Characteristics output from Stage 1, samples are classified into seven categories:
  • generic: General conversation, no specific domain
  • math_solving: Mathematical reasoning, problem-solving
  • reasoning: Logical reasoning, deduction, inference
  • code_writing: Programming, code generation/explanation
  • complex_instruction_following: Multi-step instructions, constraints
  • open_qa: Open-domain question answering, factual queries
  • closed_qa: Closed-domain QA, specific context provided
Classification uses simple rules (presence of code blocks $\rightarrow$ code_writing, math symbols $\rightarrow$ math_solving, etc.) with GPT-4o fallback for ambiguous cases.
Evaluation Dimensions: Each category evaluator assesses five dimensions on 1-5 scale:
  • Instruction Following (IF): Adherence to constraints, completeness of response relative to request
  • Contextual Alignment (CA): Relevance to conversation history, appropriate continuations
  • Accuracy (AC): Factual correctness, logical validity, absence of hallucinations
  • Completeness (CO): Thoroughness, addressing all aspects of query
  • Linguistic Clarity (LC): Grammar, coherence, readability, fluency
Category-Specific Prompts: Each category uses specialized evaluation prompt optimized for relevant quality aspects. For example:
  • math_solving: Emphasizes step-by-step correctness, formula accuracy, numerical precision
  • code_writing: Focuses on syntax validity, logic correctness, edge case handling, efficiency
  • reasoning: Evaluates logical coherence, premise-conclusion validity, absence of fallacies
Prompts include few-shot examples for calibration (2-3 examples per category).
Output Format: Evaluators return structured JSON:
"instruction_following": 4,
"explanation_IF": "Response addresses all...",
"contextual_alignment": 5,
"explanation_CA": "Perfectly aligned with...",
"accuracy": 4,
"explanation_AC": "Factually correct but...",
"completeness": 4,
"explanation_CO": "Covers main points...",
"linguistic_clarity": 5,
"explanation_LC": "Clear, well-structured..."
JSON enforcement via model’s structured output mode (GPT-4o, Claude 3.5) or constrained decoding (vLLM with guidance). Explanations provide interpretability and debugging support.
4.3 Metadata Schema and Integration
Quality metadata follows hierarchical structure compatible with OASST format:
"conversation": [...], // original dialogue
"metadata": {
"quality_characteristics": {
"heuristic_based": {
"lexical_richness": {
"ttr_score": 0.67,
"mtld_score": 89.3
"perplexity": {"score": 234.5},
"language": {
"detected": "en",
"confidence": 0.98
"conversation_stats": {
"turn_count": 4,
"avg_turn_length": 127
"LLM_based": {
"category": "math_solving",
"instruction_following": 4,
"contextual_alignment": 5,
"accuracy": 4,
"completeness": 4,
"linguistic_clarity": 5,
"explanations": {...}
This format integrates seamlessly with training pipelines:
  • Filtering: Reject samples where any dimension << threshold (e.g., accuracy << 3 for factual datasets)
  • Stratified Sampling: Balance quality distributions (e.g., equal numbers of scores 3, 4, 5)
  • Reward Modeling: Use dimension scores as auxiliary supervision signals
  • Curriculum Learning: Order training samples by difficulty (ascending average score)
  • Weighted Sampling: Sample probability proportional to quality score during training
OASST compatibility enables direct use with HuggingFace Transformers’ SFT and DPO trainers without format conversion.
@@ 6 Availability
SyGra is released as an open-source Python package and framework on PyPI111pip package: https://pypi.org/project/sygra/ (pip install sygra) and GitHub [^2] https://github.com/ServiceNow/SyGra.
@@ 7 Conclusion
Current limitations include multimodal outputs being restricted to GPT-based endpoints (other backends remain text-only), independent node operation without cross-sample reasoning, and basic agent support.
@@
Acknowledgements
We gratefully acknowledge the following people for their contributions: Nirali Popat, Sidharthenee Nayak, Nandhakumar Kandasamy, Sravan Ramachandran, Segan Subramanian, Masoud Hashemi and Rishabh Maheshwary.
@@ A.1.1 Input Sources
This configuration illustrated below LABEL:lst:data-config represents:
@@
An example configuration using a HuggingFace dataset as source and applying field renaming transformation is shown below.
data_config:
source:
# Example 1: HuggingFace dataset source
type: "hf" # HuggingFace dataset
repo_id: "google-research-datasets/mbpp" # HuggingFace repository ID
config_name: "sanitized" # Dataset configuration name
split: ["train", "validation", "prompt"] # Dataset splits to use
# OR
# Example 2: Local file source
type: "disk" # Local file source
file_path: "/path/to/data.json" # Path to input file
file_format: "json" # Format (json, jsonl, csv, parquet)
encoding: "utf-8" # File encoding
# Optional transformations to apply to the input data
transformations:
- transform: sygra.processors.data_transform.RenameFieldsTransform # Path to transformation class
params: # Parameters for the transformation
mapping:
task_id: id # Rename ’task_id’ field to ’id’
overwrite: false # Don’t overwrite existing fields
# Optional sink configuration for where to store output data
sink:
# Example 1: HuggingFace dataset sink
type: "hf" # HuggingFace dataset
repo_id: "output-dataset/synthetic-mbpp" # Where to upload the data
split: "train" # Split to write to
private: true # Create a private dataset
# OR
# Example 2: Local file sink
type: "json" # File format (json, jsonl, csv, parquet)
file_path: "/path/to/output/file.json" # Path to save the file
encoding: "utf-8" # File encoding
@@ A.1.2 Transformations
Example below shows a sample usage where the fields page, llm_extract, and type are renamed to id, text, and text_format, respectively.
Example usage of RenameFieldsTransform in YAML configura- tion. This renames selected fields to align with graph input expectations.
- transform: sygra.processors.data_transform.RenameFieldsTransform
params:
mapping:
page: id
llm_extract: text
type: text_format
@@
This transformation combines multiple records to form richer contextual input. It can skip from the beginning or end of the dataset, define how many records to combine, and how to shift the combination window. As shown below, the configuration merges two records, joining multiple fields with newline delimiters or preserving the first record’s values.
- transform: sygra.processors.data_transform.CombineRecords
params:
skip:
from_beginning: 10
from_end: 10
combine: 2
shift: 1
join_column:
page: "\$1-\$2"
[…]
type: "\$1"
model: "\$1"
metadata: "\$1"
@@
It presents a simpler configuration to exclude records from the dataset, either from the start or end. This is especially useful for filtering noisy, incomplete, or structurally incompatible entries prior to processing.
- transform: sygra.processors.data_transform.SkipRecords
params:
skip_type: "count"
count:
from_start: 10
from_end: 10
@@ A.1.3 Data Less Mode
The below YAML shows a minimal configuration that defines only an output sink.
data_config:
# No source configuration
# Only sink configuration
sink:
type: "json"
file_path: "output/synthetic_data.jsonl"